A railway train is travelling on a curve of 750m radius at the rate of 30km/hr,through what angle has it turned in 10 seconds.
1
Expert's answer
2012-09-21T11:12:49-0400
The distance traveled by the train in 10 seconds is

D = 30[km/h] * 10[s] = 30[km/h] * 1/360[h] = 1/12 [km].

So, the angle it has turned is

& alpha; = D/R = (1/12[km])/(0.750[km]) = 1/9 rad.